OpenStudy (anonymous):

What is the equation of the osculating circle that intersects the parabola {f(x; x^2) : x E R} at the point (0; 0)? Find an arc length parametrization of this circle.

OpenStudy (amistre64):

i recall something about a "k" ...

OpenStudy (amistre64):

\[r = \frac{1}{|k(t)|}\]where k is curvature

OpenStudy (amistre64):

\[k=\frac{y''}{(1+y'^2)^{3/2}}\]

OpenStudy (anonymous):

But would i need the equation of the circle to find K? like at least know the radius?

OpenStudy (amistre64):

k can be determined from y = x^2, and k is used to determine the radius the equation for the circle in this problem is centered at (0,r)

OpenStudy (amistre64):

y = x^2 y' = 2x y'' = 2 k = 2/(1+(2x)^2)^(3/2), at x=0; k = 2 r = 1/k = 1/2 x^2 + (y-1/2)^2 = 1/4 is the equation of the circle, but not in parametric form